OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = –5x + 2, solve for f(g(x)) when x = 3.

OpenStudy (anonymous):

I feel like I know how to solve but I'd like to be sure!

jhonyy9 (jhonyy9):

for f(g(x)) just substitute the g(x) in place of x inside f(X) OK ?

jhonyy9 (jhonyy9):

after that calculi this for x=3

OpenStudy (anonymous):

So f(x)=2(-5x+2)+1?

jhonyy9 (jhonyy9):

this wan being f(g(x)) not f(x) ,ok ?

jhonyy9 (jhonyy9):

so than is right just now you need calculi it

OpenStudy (anonymous):

I'm sorry what?

OpenStudy (anonymous):

@Loser66

jhonyy9 (jhonyy9):

this is above right for f(g(x))

OpenStudy (anonymous):

I'm lost.

OpenStudy (anonymous):

@johnweldon1993 would you mind helping me out? @jhonyy9 with all due respect, sir, your english isn't too great.

OpenStudy (johnweldon1993):

f(x) = 2x + 1 g(x) = -5x + 2 f(g(x)) when x = 3 G in inside the 'f' function so we focus on that first....we can plug in the x = 3 into the g(x) equation we have g(x) = -5x + 2 so g(3) = -5(3) + 2 = -15 + 2 = -13

OpenStudy (johnweldon1993):

Now..we plug in what we got as a result from g(3)...and plug it into f(x) we received an answer of -13 from g(3)...so we plug in that -13 into f(x) wherever we see an 'x'

OpenStudy (johnweldon1993):

so f(x) = 2x + 1 f(g(3)) = 2(-13) + 1 = -26 + 1 = -25

OpenStudy (anonymous):

Ahh okay! I got -13 at some point but didn't know how to carry on. Thank you so much!

OpenStudy (johnweldon1993):

Not a problem :)

OpenStudy (johnweldon1993):

And just 1 more thing...what @jhonyy9 was doing was also correct. Notice how I upfront plugged in x = 3....what jhonny9 was doing was going through the algebra first...and then plugging in the 3 so for that...he took the fact that g(x) = -5x + 2 And he plugged THAT into f(x) wherever there was an 'x' since f(x) = 2x + 1 That made it f(g(x)) = 2(-5x + 2) + 1 = -10x + 4 + 1 = -10x + 5 Now when you plug in x = 3 you get -10(3) + 5 -30 + 5 -25 Same as we received as plugging it in first *Just wanted to clear that up :)